OpenStudy (anonymous):

What are the values of x and y in this linear system? 2x + 3y = 8 x + 2y = 12 A)x = 24, y = 8 B) x = 16, y = -20 C) x = -8, y = -16 D) x = -20, y = 16 @kelliegirl33 @L-Lawliet-L

OpenStudy (anonymous):

do you prefer elimination or substitution ?

OpenStudy (anonymous):

Either or

OpenStudy (anonymous):

lets use substitution...it is sometimes easier... x + 2y = 12 x = -2y + 12 now we sub -2y + 12 in for x in the other equation 2x + 3y = 8 2(-2y + 12) + 3y = 8 -4y + 24 + 3y = 8 -4y + 3y = 8 - 24 -y = - 16 y = 16 now sub 16 in for y x + 2y = 12 x + 2(16) = 12 x + 32 = 12 x = 12 - 32 x = - 20 answer is D

OpenStudy (anonymous):

Thanks, great help! :)